# Juntao Zhao1,*    Wenhao Lu2,†    Sheng Wang1,†    Lingpeng Kong1    Chuan Wu1

### v2

@@ 3.2   Advantages of High Acceptance Rate
A key advantage of QSpec lies in its superior acceptance rate. While state-of-the-art speculative decoding approaches such as EAGLE Li et al. (2024a) and Medusa Cai et al. (2024) rely on compressed draft models and tree-structured drafting to balance speed and accuracy, QSpec achieves high performance without such compromises. We analyze how QSpec demonstrates superior performance compared to these methods in quantized settings below.
Tree-Structured Drafting. As shown in Figure 4 (a), tree-structured drafting (Chen et al., 2024; Miao et al., 2024; Li et al., 2024a, b) is a widely adopted technique to improve acceptance rates in speculative decoding. Instead of generating a single draft token at a time using the draft model $\mathcal{M}_{d}$, $\mathcal{M}_{d}$ select $k$ (i.e., topk) tokens and infers $\gamma$ times to form a draft tree with depth $\gamma$. The target model $\mathcal{M}$ then verifies the tokens using masked tree attention, with the exact verification strategy depending on the sampling method (Cai et al., 2024; Miao et al., 2024). Under greedy sampling, the highest-probability token at each position is selected, forming the longest branch with the common prefix as the accepted sequence. In this tree, root node presents the past sequence, any token $t_{i}$ has a path from the root node $r$ to $t_{i}$, denoted by $Path(r,t_{i})$, consisting of ancestors $a_{1},\ldots,a_{j}$.
Expected average accepted token number. Let $p_{a}(t)$ be the probability of accepting a token $t$. The average accepted token number is given by:
$H(k):=\sum_{l}^{\gamma}\sum_{i=1}^{k^{l}}\prod_{j\in Path(r,t_{i})}p_{a}(a_{j}),$ (2)
@@
where $k$ controls the tree’s branching factor (width) . Larger $k$ indicates larger expected token acceptance. When $k=1$, this method reduces to standard speculative decoding.
@@
Cost Analysis. Denote by $C_{d}(\cdot)$ and $C_{p}(\cdot)$ the computation cost for draft and verify, The first argument indicates the number of sequences, and the second argument is the prefix sequence length. For a prefix with length $s$, we have per-valid-token cost $v$ by dividing the total cost by the number of accepted tokens:
@@
$v&=\frac{\overbrace{C_{d}(1;s)+\cdots+C_{d}\bigl{(}k^{\gamma{}-1};s+\gamma{}-1\bigr{)}}^{\text{Drafting cost}}\;+\;\overbrace{C_{p}\bigl{(}k^{\gamma{}-1};s+\gamma{}\bigr{)}}^{\text{Verify cost}}}{\underbrace{H(k)}_{\text{Accepted tokens}}}\sim\frac{C(k)}{H(k)}.\end{aligned}$
Batch serving efficiency of QSpec. From Equation 3, efficient tree-structured drafting benefits from less extra computational overhead $\Delta C(k)_{\text{SD}}$ against acceptance rate gains $\Delta H(k)$, as illustrated in Figure 4(c). QSpec advances this paradigm in batched serving. As shown in Figure 4(b), QSpec employs W4A4 quantization for drafting. While initially exhibiting two limitations - suboptimal hardware utilization during the drafting and higher computational costs compared to small/pruned models - these challenges diminish with increased batch sizes. Beyond critical batch size $bs^{*}$, QSpec’s draft phase transits from memory-bound to compute-bound, achieving superior FLOP/s with low-precision. Also, by eschewing penalties for acceptance rate maintenance, QSpec reduces costs by a factor of $k$. As a result, in Figure 4(c), $\Delta C(k)_{\textsc{QSpec}{}}$ curve maintains a consistently shallower slope than $\Delta C(k)_{\text{SD}}$ for $bs>bs^{*}$, enabling higher token generation throughput. This observation applies to both static (e.g., SpecInfer (Miao et al., 2024), Sequoia (Chen et al., 2024)) and dynamic (Wang et al., 2024; Li et al., 2024b) tree-structured drafting methods, all of which exhibit similar behaviors.
@@ 2. Q2:
    Does QSpec accelerate high-precision weight-only quantization methods and surpass SOTA speculative decoding method in quantized scenarios? (Sec.4.3)
@@ 3. Q3:
    How do various factors (e.g., quantization method, sequence length) influence QSpec’s acceptance rate and acceleration performance? (Sec. 4.4)
@@ 4.1   General Setup
Benchmarks. We assess QSpec with two primary criteria: (1) generation fidelity and (2) end-to-end serving speedup. For fidelity evaluation, we adopt not only traditional tasks, including PIQA (500, 10-shot) (Bisk et al., 2020), WinoGrande (500, 5-shot) (Sakaguchi et al., 2019), and WikiText2 (Merity et al., 2016), but also challenging multi-step reasoning tasks such as GSM8K (All, 8-shot) (Cobbe et al., 2021), MATH (All, 4-shot) (Hendrycks et al., 2021), MBPP (200, 0-shot) (Austin et al., 2021), and HumanEval (All, 0-shot) (Chen et al., 2021). To measure the acceleration, we use all the above reasoning tasks and two additional chatbot datasets, namely ShareGPT (RyokoAI, 2021) and LMsys-1K (Zheng et al., 2023). Due to space constraints, we present results for GSM8K, HumanEval, and LMsys-1K in the main text, with remaining results detailed in the Appendix A. Following the setup of Atom (Zhao et al., 2024b), we randomly sampled the dataset for the request prompts to reduce the workload. Due to memory limitations, we vary the batch size from 8 to 32 and serve all requests in a first-come, first-served (FCFS) manner. Once any request is finished, we refill the batch, adhering to the continuous batching approach of ORCA (Yu et al., 2022). All experiments use greedy sampling for token generation.
@@
Implementations. All experiments are performed on a node equipped with four NVIDIA L20 GPUs (48GB HBM each) running CUDA 12.5. To demonstrate the versatility of QSpec, we implement two SOTA 4-bit quantization methods, namely Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024). For W4A16 configurations, we incorporate AWQ-style (Lin et al., 2024a) weight dequantization logic for runtime inference. We select Atom to showcase the acceleration of QSpec. We use these Group-wise quantization schemes with a group size of 128. With the draft token length $\gamma$ as 3, we simulate the performance of QSpec by initially employing fake quantization to fully emulate the execution flow, encompassing both the draft and verify stages of QSpec. Subsequently, we replay the collected traces with real kernel execution to accurately reproduce the latency.
Baselines. We evaluate QSpec against baseline quantization configurations: W4A4, W4A16, and W16A16. For comparison with SOTA speculative decoding approaches, we include EAGLE (Li et al., 2024a), which employs a pruned draft model with tree-structured speculative decoding. In our quantized experiments, we utilize FP16 precision for the EAGLE draft model and EAGLE-Quant (W4A16) for the target model. This choice was necessitated by two factors: (1) the official EAGLE quantization implementation (fast-gpt) lacks efficient batching support, and (2) applying GPTQ quantization to the EAGLE draft model resulted in substantial degradation of the acceptance rate.
4.2   Fidelity Evaluation
@@
QSpec effectively maintains the generation quality of W4A16, whereas W4A4 does not. As listed in Table 3, with the draft verification of W4A16, QSpec exhibits only minimal performance fluctuations compared to W4A16. This negligible variation may stem from the nondeterministic algorithms of PyTorch (PyTorch Contributors, 2024) or occasional cases where two tokens have the same maximum prediction probability. In contrast, W4A4 experiences a substantial performance decline exceeding 10% across most tasks, with the reduction becoming more pronounced as task difficulty increases. For instance, compared to GSM8K and MBPP, the performance drop for W4A4 is much greater on the more challenging MATH and HumanEval tasks, showing declines of 51.11% and 38.73%, respectively. On the other hand, this also highlights the higher sensitivity of multi-step reasoning tasks to the negative effects of quantization compared to regular tasks, such as WikiText-2and WinoGrande. This observation aligns with our earlier analysis in Sec. 2, encouraging incorporating multi-step reasoning tasks into quantization evaluation.
4.3   Acceleration Evaluation
Table 4: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. ”Avg.” denotes the average speedup ratio for the corresponding row or column. $\dagger$” indicates the failure of W4A16 kernels to support the configuration. W16A16 results (when OOM) are collected using pipeline parallelism Harlap et al. (2018) with multiple cards.
@@ 4.3   Acceleration Evaluation
QSpec 8 594.1 (1.41$\times$) 723.6 (1.35$\times$) 738.8 (1.32$\times$)
16 811.5 (1.40$\times$) 1042.1 (1.30$\times$) 1171.4 (1.27$\times$)
32 1030.4 (1.42$\times$) 1248.5 (1.31$\times$) 1576.0 (1.18$\times$)
Avg. 1.41$\times$ 1.32$\times$ 1.25$\times$
7B W16A16 8 213.4 316.7 285.3
16 290.3 505.1 441.6
32 340.9 663.6 564.2
W4A4 8 349.5 471.2 419.4
@@
QSpec 8 253.7 (1.54$\times$) 350.9 (1.46$\times$) 310.3 (1.41$\times$)
16 359.8 (1.55$\times$) 555.2 (1.36$\times$) 473.1 (1.32$\times$)
32 441.8 (1.64$\times$) 749.4 (1.35$\times$) 628.4 (1.34$\times$)
Avg. 1.58$\times$ 1.39$\times$ 1.36$\times$
13B W16A16 8 121.9 182.0 160.1
16 169.6 291.0 243.0
32 202.4 423.5 334.2
W4A4 8 194.7 261.5 228.2
@@
32 $\dagger$ 365.5 287.4
QSpec 8 148.2 (1.56$\times$) 201.2 (1.44$\times$) 174.0 (1.36$\times$)
16 212.8 (1.56$\times$) 323.3 (1.36$\times$) 266.9 (1.29$\times$)
32 266.6 ($\dagger$) 483.0 (1.32$\times$) 379.3 (1.32$\times$)
Avg. 1.56$\times$ 1.37$\times$ 1.32$\times$
QSpec exhibits a substantial efficiency boost compared to W4A16. In Table 4, we present the token generation throughput for both QSpec and W4A16 across different model sizes, quantization configurations, and batch sizes on diverse datasets. On average, QSpec achieves a throughput increase of 1.38×1.38$\times$1.38 × over W4A16 across all settings, with a peak improvement of 1.64×1.64$\times$1.64 ×.
Larger models tend to yield better speedup ratios. We find a gradually-increasing acceleration as the base model scales up. This overall upward trend indicates a promising outlook for our approach with larger models, although further experiments are needed for confirmation. Due to resource limitations, this will be addressed in future work.
Latency composition. As illustrated in Figure 5, we compute the per-valid-token latency by dividing the total latency by only the number of accepted tokens before averaging on all evaluation datasets. Notably, QSpec achieves remarkable latency savings ranging from 28.5% to 39.7%. Besides, the per-token latency is further decomposed into two parts: draft and verify. Clearly, the primary gains of QSpec arise from the rapid drafting capability and the reduced latency achieved through the parallel verification of multiple tokens.
@@ 16 OOM OOM OOM
Memory and batching efficiency over speculative decoding. Table 5 compares QSpec and EAGLE on Llama-2-7b-chat-hf. EAGLE delivers optimal performance for single-sequence inputs (batch size=1). However, its efficiency degrades as the batch size grows (8 and 16). This observation aligns with our earlier analysis in Sec. 3.2: EAGLE’s tree-structured drafting mechanism, designed to reconcile discrepancies between the draft and target models, introduces additional latency that reduces the gains from higher acceptance rates in batched serving under quantization. Furthermore, the increasing key-value (KV) storage of EAGLE’s draft model leads to out-of-memory (OOM) issues at batch size 16. In contrast, QSPEC demonstrates superior scalability and memory efficiency.
@@ 4.4   Ablation Study
Ablation on draft token length. To assess parameter sensitivity, we vary the draft token lengths $\gamma$, the sole hyper-parameter of QSpec, from 2 to 6 across all the benchmarks using Llama3.2-3b and Llama3-8b-instruct models. As depicted in Figure 6, an increase in $\gamma$ leads to a gradual decline in the token acceptance rate, since all subsequent tokens are discarded once a token is rejected. Nevertheless, even at $\gamma=6$, the token acceptance rate remains relatively high, approximately $74\%$, compared to $28\sim 58\%$ in 160m-7b draft-target model pair under $\gamma=5$ in conventional speculative decoding (Liu et al., 2024). Additionally, a consistent improvement in throughput is observed compared to W4A16, indicating the robustness of QSpec with respect to $\gamma$. More comprehensive comparison is provided in Appendix A.2.
Quantization Method ShareGPT MATH (4-shot) MBPP (0-shot)
Atom 83.8 89.4 88.6
QuaRot 81.6 88.9 85.4
Ablation on base quantization method. As shown in Table 6, QSpec consistently achieves high acceptance rates across diverse quantization methods and datasets.
@@ 5   Related work
Speculative Decoding leverages a draft model to generate candidate tokens, which are then validated by a target model (Leviathan et al., 2023). Recent research has primarily focused on improving the acceptance rate and generation speed of candidate tokens. SpecInfer (Miao et al., 2024) introduces a boost-tuned small language model to generate candidate tokens in tree structures, enabling single-pass verification. In contrast, EAGLE (Li et al., 2024a) adopts an aggressive pruning strategy for the draft model’s architecture, allowing penultimate layer feature prediction with minimal computational overhead. Self-speculative decoding, a subset of this technique, employs a single model for both draft generation and verification. LayerSkip (Elhoushi et al., 2024) introduces a training methodology for early exit with layer drop, subsequently verifying partially generated tokens through full model inference. Medusa (Cai et al., 2024) augments the original LLM with additional heads atop the final hidden state while relaxing the acceptance policy. However, these approaches inevitably require retraining of the original model, which can be computationally expensive and time-consuming. We further demonstrate their deficiency in batched serving under quantization scenario.
@@ 6   Conclusion
In this paper, we begin by validating that multi-step reasoning tasks can capture performance degradation incurred by activation quantization more sensitively and consistently than current evaluation protocols, advocating for their incorporation for a more comprehensive assessment. With observed high token-level similarities, we introduce QSpec, a novel quantization paradigm that seamlessly synergizes two complementary weight-shared quantization schemes with speculative decoding. Empirically, QSpec achieves up remarkable acceleration without any quality sacrifice across diverse settings and baselines. Alongside consistent memory consumption and a plug-and-play property, these advantages distinguish QSpec from any existing solution, promising it for high-fidelity quantization deployment, particularly in memory-constrained scenarios.
7   Impact Statement

### v3

@@ 3.2 Advantages of High Acceptance Rate
A key advantage of QSpec lies in its superior acceptance rate. While state-of-the-art speculative decoding approaches such as EAGLE Li et al. (2024b) and Medusa Cai et al. (2024) rely on distilled draft models and tree-structured drafting to balance speed and accuracy, QSpec achieves high performance without such compromises. We analyze how QSpec demonstrates superior performance compared to these methods in quantized settings below.
@@
Expected average accepted token number. Let $p_{a}(t)$ be the probability of accepting a token $t$. The average accepted token number is given by Equation 2, where $k$ controls the tree’s branching factor (width) . Larger $k$ indicates larger expected token acceptance. When $k=1$, this method reduces to standard speculative decoding.
$H(k):=\sum_{l}^{\gamma}\sum_{i=1}^{k^{l}}\prod_{j\in Path(r,t_{i})}p_{a}(a_{j}).$ (2)
@@
Cost Analysis. Denote by $C_{d}(\cdot)$ and $C_{p}(\cdot)$ the computation cost for draft and verify, The first argument indicates the number of sequences in a draft tree, and the second argument is the prefix sequence length. For a prefix with length $s$, we have per-valid-token cost $v$ by dividing the total cost by the number of accepted tokens in Equation:
@@
$\begin{aligned} v&=\frac{\overbrace{C_{d}(1;s)+\cdots+C_{d}\bigl(k^{\gamma{}-1};s+\gamma{}-1\bigr)}^{\text{Drafting cost}}\;+\;\overbrace{C_{p}\bigl(k^{\gamma{}-1};s+\gamma{}\bigr)}^{\text{Verify cost}}}{\underbrace{H(k)}_{\text{Accepted tokens}}}\sim\frac{C(k)}{H(k)}.\end{aligned}$
Batch serving efficiency of QSpec. The efficacy of speculative decoding relies on the assumption that verification cost $C_{p}$ approximates the target model’s decoding cost $C_{d}$, leveraging underutilized resources for parallel verification.(Chen et al., 2023; Leviathan et al., 2023; Yan et al., 2025) This holds in single-request, memory-bound scenarios. However, tree-structured drafting, with verification cost $C_{p}(k^{\gamma-1};s+\gamma)$, scales poorly as $k^{\gamma-1}$ far exceeds non-tree sequence length $\gamma$. Therefore, in batched serving, tree-based $C_{p}$ enters compute-bound territory easily, significantly surpassing $C_{d}$ and incurring high memory overhead, invalidating the assumption. QSpec, using W4A4 quantization, avoids tree structures, maintaining low $C_{p}$ and memory usage and balancing the loads between drafting and verification stages, QSpec delays the onset of this inefficiency. In Section 4, QSpec demonstrates excellent speedup even at a batch size of 32, also outperforming EAGLE (Li et al., 2024b) in comparison experiments.
@@ 2. Q2:
    Does QSpec speed up high-precision weight-only quantization methods, and surpass speculative decoding methods in quantized scenarios? (Sec.4.3)
@@ 3. Q3:
    How do various factors (e.g., quantization method, sequence length) influence the acceptance rate and acceleration performance of QSpec? (Sec. 4.4)
@@ 4.1 General Setup
Benchmarks. We assess QSpec with two primary criteria: (1) generation fidelity and (2) end-to-end serving speedup. For fidelity evaluation, we adopt not only traditional tasks, including PIQA (500, 10-shot) (Bisk et al., 2020), WinoGrande (500, 5-shot) (Sakaguchi et al., 2019), and WikiText2 (Merity et al., 2016), but also challenging multi-step reasoning tasks such as GSM8K (All, 8-shot) (Cobbe et al., 2021), MATH (All, 4-shot) (Hendrycks et al., 2021), MBPP (200, 0-shot) (Austin et al., 2021), and HumanEval (All, 0-shot) (Chen et al., 2021). To measure the acceleration, we use all the above reasoning tasks and two additional chatbot datasets, namely ShareGPT (RyokoAI, 2021) and LMsys-1K (Zheng et al., 2023a). Due to space constraints, we present results for GSM8K, HumanEval, and LMsys-1K in the main text, with remaining results detailed in the Appendix A. Following the setup of Atom (Zhao et al., 2024b), we randomly sampled the dataset for the request prompts to reduce the workload. Due to memory limitations, we vary the batch size from 8 to 32 and serve all requests in a first-come, first-served (FCFS) manner. Once any request is finished, we refill the batch, adhering to the continuous batching approach of ORCA (Yu et al., 2022). All experiments use greedy sampling for token generation.
@@
Implementations. All experiments are performed on a node equipped with four NVIDIA L20 GPUs (48GB HBM each) running CUDA 12.5. To demonstrate the versatility of QSpec, we implement two SOTA 4-bit quantization methods, namely Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024). For W4A16 configurations, we incorporate AWQ-style (Lin et al., 2024a) weight dequantization logic for runtime inference. We select Atom to showcase the acceleration of QSpec. We use these Group-wise quantization schemes with a group size of 128. We configure the default draft token length $\gamma$ to 3. The implementation of QSpec follows Atom’s setup, incorporating flashinfer (Ye et al., 2025).
Baselines. We evaluate QSpec against baseline quantization configurations: W4A4, W4A16, and W16A16. We also include EAGLE (Li et al., 2024b), which employs a pruned draft model with tree-structured speculative decoding, as a baseline to compare with regular speculative decoding. In our quantized experiments, we utilize FP16 precision for the EAGLE draft model and EAGLE-Quant (W4A16) for the target model. This choice was necessitated by two factors: (1) the official EAGLE quantization implementation (fast-gpt) lacks efficient batching support, and (2) applying GPTQ quantization to the EAGLE draft model resulted in substantial degradation of the acceptance rate.
4.2 Fidelity Evaluation
QSpec effectively maintains the generation quality of W4A16, whereas W4A4 does not. As listed in Table 3, with the draft verification of W4A16, QSpec exhibits only minimal performance fluctuations compared to W4A16. This negligible variation may stem from the non-deterministic algorithms of PyTorch (PyTorch Contributors, 2024) or occasional cases where two tokens have the same maximum prediction probability. In contrast, W4A4 experiences a substantial performance decline exceeding 10% across most tasks, with the reduction becoming more pronounced as task difficulty increases. For instance, compared to GSM8K and MBPP, the performance drop for W4A4 is much greater on the more challenging MATH and HumanEval tasks, showing declines of 51.11% and 38.73%, respectively. On the other hand, this also highlights the higher sensitivity of multi-step reasoning tasks to the negative effects of quantization compared to regular tasks, such as WikiText-2and WinoGrande. This observation fully aligns with our earlier analysis in Sec. 2, encouraging the incorporation of multi-step reasoning tasks into quantization evaluation.
Table 4: Comparison of token generation throughput across different model sizes, quantization configurations, and batch sizes for various datasets. All values are measured in token/s. "Avg." denotes the average speedup ratio for the corresponding row or column.
@@ 4.2 Fidelity Evaluation
[…]
7B W4A4 8 349.5 471.2 419.4
@@
[…]
13B W4A4 8 194.7 261.5 228.2
@@
32 207.5 365.5 287.4
[…]
4.3 Acceleration Evaluation
QSpec exhibits a substantial efficiency boost compared to W4A16. In Table 4, we present the token generation throughput for both QSpec and W4A16 across different model sizes, quantization configurations, and batch sizes on diverse datasets. On average, QSpec achieves a throughput increase of $1.38\times$ over W4A16 across all settings, with a peak improvement of $1.64\times$.
Larger models tend to yield better speedup ratios. We observe a consistent acceleration trend as the base model scales, demonstrating the promising scalability of our approach with larger models. While further validation is needed, resource constraints necessitate addressing this in future work.
QSpec reduces latency through fast drafting and parallel verifying. As illustrated in Figure 4, we compute the per-valid-token latency by dividing the total latency by only the number of accepted tokens before averaging on all evaluation datasets. Notably, QSpec achieves remarkable latency savings ranging from 26.5% to 30.6%. Besides, the per-token latency is further decomposed into two parts: draft and verify. Clearly, the primary gains of QSpec arise from the rapid drafting capability and the reduced latency achieved through the parallel verification of multiple tokens.
@@ 16 OOM OOM OOM
QSpec offers better memory and batching efficiency than prior speculative decoding methods. Table 5 compares QSpec and EAGLE on Llama-2-7b-chat-hf. EAGLE delivers optimal performance for single-sequence inputs (batch size=1). However, its efficiency degrades as the batch size grows (8 and 16). This observation aligns with our earlier analysis in Sec. 3.2: EAGLE’s tree-structured drafting mechanism, designed to reconcile discrepancies between the draft and target models, introduces additional latency, and reduces the gains from higher acceptance rates in batched serving under quantization. Furthermore, the increasing key-value (KV) storage of EAGLE’s draft model leads to out-of-memory (OOM) issues at batch size 16. In contrast, QSpec demonstrates superior scalability and memory efficiency.
[…]
@@ 4.4 Ablation Study
Ablation on draft token length. To assess parameter sensitivity, we vary the draft token lengths $\gamma$, the only hyper-parameter of QSpec, from 2 to 6 across all the benchmarks using Llama3.2-3b and Llama3-8b-instruct models. As depicted in Figure 5, an increase in $\gamma$ leads to a gradual decline in the token acceptance rate, since all subsequent tokens are discarded once a token is rejected. Nevertheless, even at $\gamma=6$, the token acceptance rate remains relatively high, approximately $74\%$, compared to $28\sim 58\%$ in 160m-7b draft-target model pair under $\gamma=5$ in conventional speculative decoding (Liu et al., 2024). Additionally, a consistent improvement in throughput is observed compared to W4A16, indicating the robustness of QSpec with respect to $\gamma$. More comprehensive comparison is provided in Appendix A.2.
@@
Ablation on base quantization method. As shown in Appendix Table 9, QSpec consistently achieves high acceptance rates across diverse quantization methods and datasets.
@@ 5 Related Work
This work builds on two lines of research: quantization and speculative decoding. Weight-only quantization (W4A16) offers better accuracy, while joint weight-activation methods (W4A4) enable faster inference but degrade performance on complex tasks. Speculative decoding improves efficiency by verifying drafted tokens, but existing approaches typically require retraining and are less effective under quantization. We provide detailed comparisons in Appendix D.
@@ 6 Conclusion
We begin by showing that multi-step reasoning tasks reveal performance degradation from activation quantization more clearly than current evaluation protocols, and encourage their incorporation for more comprehensive assessment. Leveraging high token-level similarities, we propose QSpec, a novel quantization paradigm that seamlessly combines two complementary weight-shared quantization schemes through speculative decoding. Empirically, QSpec delivers significant acceleration—up to $1.64\times$—without quality loss across diverse settings. With consistent memory usage and a plug-and-play design, QSpec offers a practical and scalable solution for high-fidelity quantization, especially under memory constraints.
Acknowledgements
This work was supported in part by grants from Hong Kong RGC under the contracts 17204423, 17205824, C7004-22G (CRF), C5032-23G (CRF) and T43-513/23-N (TRS).
7 Limitations and Impacts
7.1 Impact Statement